OpenStudy (anonymous):

how many solutions are there to the equation, 6x+35+9x=15(x+4)-25?

OpenStudy (anonymous):

\[6x+35+9x=15x+35\] \[15(x+4)-25=15x+60-25=15x+35\] now we have 15x+35=15x+35 these equations are equal. there is no limit to the number of solutions, it is infinite. any number you put in for x on one side will yield the same result on the other. take x=2 for example; 6(2)+35+9(2)=65 15(2+4)-25=65 equal. how about x=-3 6(-3)+35+9(-3)=-10 15(-3+4)-25=-10 equal again.

OpenStudy (turingtest):

Infinite